MILLIMETER WAVE SPECTROSCOPY OF THE VINYL RADICAL (HDC=CH) GENERATED BY UV LASER PHOTOLYSIS IN A PULSED SUPERSONIC JET EXPANSION

\underline{MASATO HAYASHI}, KENSUKE HARADA and KEIICHI TANAKA, Department of Chemistry, Faculty of Science, Kyushu University 33, Hakozaki, Higashiku, Fukuoka 812-8581, Japan.

The vinyl radical has tunneling motion of acetylenic (CH) proton. The proton tunneling transitions of the vinyl radical, both for the normal (H_{2}CCH) and deuterated species (H_{2}CCD), were observed by millimeter wave spectroscopy and the proton tunneling splittings of H_{2}CCH and H_{2}CCD in the ground state have been determined 16271.8429(59) MHz and 1164.861(20) MHz, respectively. The HDCCH is a mono-deuteride of the methylenic (CH_{2}) proton of the vinyl radical, and may have two local potential minima; cis- and trans-forms, where cis-HDCCH has an unpaired electron on the same side of deuterium with respect to the C=C double bond, while trans-HDCCH on the opposite side. If the energy difference Δ_{ct} between the ground state of trans-HDCCH and that of cis-HDCCH is small enough, the proton tunneling transition will be observed. On the other hand, if Δ_{ct} is large (Δ_{ct} ≥ 1 cm^{−1}), the b-type rotational transitions of ΔK_{a} = ±1 will be observed in cis- and trans-tautomers instead of the tunneling transition between the cis- and trans-forms. In the present study, the rotational spectra of the cis-HDCCH have been observed by millimeter wave spectroscopy combined with a pulsed supersonic jet technique. The HDCCH radical was generated by 193 nm excimer laser photolysis of mono-deuteride vinyl chloride(HDCCHCl). The a-type rotational transitions of the cis-HDCCH,N_{K_{a}K_{c}}=2_{02}−1_{01}, 3_{03}−2_{02}, were observed in the frequency region of 109−164 GHz. The observed rotational lines were split into fine and hyperfine components due to spin-rotation interaction and spin-nuclear spin interaction. The molecular constants such as rotational constants, spin-rotation interaction constant, and hyperfine interaction constants were determined by least squares fitting. These molecular constants are consistent with the results of the millimeter wave spectroscopy of H_{2}CCH and H_{2}CCD in gas phase, and result of ESR study in Ar matrix ^{a}. The rotational spectra of the trans-HDCCH have not been observed in supersonic jet (15 K). This fact suggests that the ground state of trans-HDCCH is much higher than that of cis-HDCCH. The zero point energy difference Δ_{ct} between cis- and trans-forms, calculated to be 38 cm^{−1} from the CCSD(T)/cc-pVTZ calculation, supports the present result by millimeter wave spectroscopy.

aP. H. Kasai,J. Am. Chem. Soc. 94, 5950 (1972)